The reaction of magnesium with nitrogen produces magnesium nitride, as follows. 3 Mg(s) + N2(g) → Mg3N2(s) If the reaction is started with 2.24 mol Mg and 0.848 mol N2, find the following.

(a) the limiting reactant

(b) the excess reactant

(c) The number of moles of magnesium nitride produced
Answer #1

as per reaction 2.24 mol Mg requires ( 1/3) x 2.24 = 0.7467 moles N2 but we had 0.848 mol N2

a) Limiting reactant is Mg

b) Excess reactanat is N2

c) Mg3N2 moles = ( 1/3) Mg moles = ( 1/3) x 2.24 = 0.7467